# Lexicon3D: Probing Visual Foundation Models for Complex 3D Scene Understanding

### v1

@@ 3.1 A Unified Probing Framework
Table 2: Evaluation of vision-language reasoning on ScanQA [5] and SQA3D [52] datasets. Top-2 results for each metric are shown in red and green, respectively. Prior method 3D-LLM results are shown for reference, indicating relative position of our evaluation results with respect to leading models trained on this task.
Figure 4: Evaluation curves on the ScanQA benchmark. DINOv2 exhibits clear superior performance.
3.2 Vision-Language Reasoning
The vision-language reasoning task requires a model to engage in dialogues or answer questions about global understanding and local concepts and objects related to a given complex 3D indoor scene. Following [35, 108], we formulate this as a visual-question answering (VQA) task using Large Language Models (LLMs) as the backbone – given a 3D scene from multi-view images and point clouds, and a user-prompt question, the LLMs are asked to generate the answer to the question in an auto-regressive way. This task encompasses universal language-guided reasoning of the complex indoor scene, ranging from global layout to local details.
Datasets and optimization.   We evaluate the performance on two challenging indoor 3D VQA datasets: ScanQA [5] and SQA3D [52]. Following the evaluation methodology of [35, 52, 5, 55], we report the metrics BLEU [61], ROUGE [47], METEOR [7], and CIDEr [82]. We finetune a Q-Former module [46] to align features from different encoders to the LLM input space. More datasets and optimization details are provided in the supplementary material.
Evaluation results.   Table 2 and Figure 4 present the results of our evaluation. We observe that image and video encoders generally outperform the 3D point encoder, with DINOv2 achieving the best performance, followed closely by V-JEPA and SVD. Interestingly, we find that for LSeg and CLIP, which are pretrained by language guidance, their advantage in language alignment does not translate into superior performance on the LLM-guided VQA task. This finding challenges the common practice of using language-pretrained visual foundation models [65, 45, 46, 44] as default encoders for LLM-based vision-language reasoning tasks. Instead, it suggests the importance of considering a wider range of encoders, such as DINOv2 and V-JEPA, to support such models.
@@ 3.3 Visual Grounding
Datasets.   We evaluate on the ScanRefer [16] dataset, which provides 51K text descriptions of 11K objects in 800 ScanNet scenes [20]. We report accuracy for unique, multiple, and overall categories, with unique referring to instances that have a unique semantic class in a given scene (easier).
@@ 3.4 Semantic Segmentation
Table 4: Results of semantic segmentation.
@@ 3.5 Registration: Geometric Correspondence
To evaluate the geometric information contained in the foundation model features, we design the following new task, partial scene registration, based on the point cloud registration [95, 49] task. From a complete point cloud representing the entire scene, we sample a pair of point clouds $P_{1}\in\mathbb{R}^{N_{1}\times 3}$ and $P_{2}\in\mathbb{R}^{N_{2}\times 3}$ within the scene, where $P_{1}$ and $P_{2}$ contain all the points that can be observed in two sets of consecutive views, respectively. Our goal is to find the homography matrix $H$ that correctly transforms the points in $P_{1}$ to register with $P_{2}$. Compared to the semantic segmentation task evaluated in Section 3.4, the partial scene registration task requires the foundation model features to have the capability of finding geometric correspondence for registration, which cannot be achieved simply by finding the correspondence according to semantic understanding. For example, in semantic correspondence, we may find two semantically similar points, one on the left side of the sofa in $P_{1}$, while the other on the right side of the sofa in $P_{2}$. As a result, if we register the two partial point clouds solely based on semantic correspondence, we will fail to find the correct homography to align one point cloud with the other. The foundation model features need to be equipped with geometric understanding capability to achieve decent performance on our partial scene registration task.
Dataset and probe head. We build our partial scene registration benchmark based on the ScanNet [20] dataset. For each scene in ScanNet, we choose views #0 $\sim$ #31 and views #32 $\sim$ #63 for rendering $P_{1}$ and $P_{2}$, respectively, so that they can have a certain level of overlap that allows the registration of two partial point clouds. Afterwards, $P_{2}$ is transformed by a homography $H$ that consists of a rotation $\mathbf{R}\in\mathrm{SO}(3)$ and a translation $\mathbf{t}\in\mathbb{R}^{3}$. $\mathbf{R}$ is created by a randomly generated quaternion $\mathbf{q}\in\mathbb{R}^{4}$ for each scene, while each component of $\mathbf{t}$ is randomly sampled from the uniform distribution $[-1.0\mathrm{m},1.0\mathrm{m}]$. We follow REGTR [95] to adopt a transformer cross-encoder module, followed by a lightweight decoder, to obtain the corresponding position of each point in the other point cloud. More details on dataset and optimization are provided in the supplementary material.
Model RR@0.05m (%) RR@0.1m (%) RR@0.2m (%) RRE () RTE (m)
@@ 4 Analysis
The purpose of this section is to provide additional exploration towards the optimal usage of visual foundation models. The selection of encoding methods requires consideration of the trade-off between memory usage, running time, and performance. We will dive into complexity analysis and the study of design choices for various and a combination of foundation models.
@@ 4.1 Complexity Analysis
We compare memory usage, computation time, and model performance (vision-language reasoning on ScanQA) in Table 6 and Figure 6. Our findings show that image encoders generally require less time to process a sample compared to video and 3D encoders. And diffusion-based models, when used for feature extraction, require significantly more memory than other discriminative models. Noticeably, the drawbacks in running time become evident for 2D backbones, especially image encoders, when attempting to obtain a scene embedding by aggregating multi-view image embeddings. To illustrate this, we consider a 300-frame video as an exemplar of posed 2D information for a complex scene (a 10-second video at 30 FPS). As the length of the video increases, 2D methods, which necessitate feature extraction for each image frame, rapidly consume a substantial amount of time to process a single scene. In contrast, a 3D point encoder requires significantly less time to process a scene. Nevertheless, 3D encoders exhibit relatively poor model performance, which can be attributed to the scarcity of training data. To fully demonstrate their potential in scene understanding tasks, efforts should be directed toward enhancing the generalizability of 3D foundation models. All analyses and computations were conducted on an Nvidia A100 GPU.
@@ 4.2 Ablation Study – Insights into Optimal Usage of Visual Foundation Models
Combination of multiple encoders.   We explore whether a mixture of foundation models (experts) has the potential to strengthen the capability of 3D scene understanding. We experiment on the 3D semantic segmentation task with three feature sources: LSeg, StableDiffusion, and Swin3D. When combining different feature sources, we concatenate all features along the channel dimension for every point in the point cloud. The results are shown in Figure 8. After combining features from different sources, there exists potential that the semantic understanding capability can be boosted in a mixture of experts manner. However, it is not necessarily true that combining the best features will lead to the best performance. For example, LSeg (1) has stronger capability on semantic segmentation than StableDiffusion (2) and Swin3D (3) individually, but it is StableDiffusion + Swin3D (2+3) that reaches the best performance when combining two features together.
Appendix (supplementary material).   The appendix offers a comprehensive introduction of all of our evaluated models and additional experiment details, and includes more visualization and ablation experiments. We also elaborate on the limitations, broader impact, and future direction of our work.
5 Conclusion
This paper presents the first comprehensive analysis of leveraging visual foundation models for complex 3D scene understanding. We explore the strengths and weaknesses of models designed for various modalities and trained with different objectives. Our study reveals the superior performance of DINOv2, the advantages of video models in object-level tasks, and the benefits of diffusion models in geometric registration tasks. Surprisingly, we find limitations of language-pretrained models in language-related tasks. The extensive analysis suggests that a more flexible encoder selection can play a crucial role in future scene understanding and multi-modal reasoning tasks.
Supplementary Material
Appendix A Additional Experiment Details
In this section, we provide a detailed introduction of all the visual foundation models we have evaluated, including the checkpoints we use and how we extract feature representations from the encoder backbones. Our code will be publicly released for replication and further evaluation.
A.1 Evaluated Visual Foundation Models
Our evaluation and analysis are conducted mainly on the seven models listed in Table 1 in the main body paper. We have chosen models such that they cover most of the backbones used by recent 3D scene understanding and reasoning work. In this part, we discuss all the models we have used in our experiments and explain their pretraining object, the dataset used for pretraining, the public checkpoints we choose, and the method we leverage to extract features from their backbones. We start with image foundation models, and then video and 3D models.
DINOv2   [58]. DINOv2 leverages an image-wise contrastive objective by minimizing the distance of features from the same samples, and maximizing those from different samples. They also include a patch-wise denoising objective by performing reconstruction from masked inputs. They train their model on a large-scale image dataset, LVD-142M [58], which contains 142 million unlabeled images. We take the standard DINOv2 implementation  [^1] https://github.com/facebookresearch/dinov2 and use the pretrained ViT-L/14 checkpoint for our evaluations.
LSeg   [44]. LSeg aims to align visual features from images with corresponding semantic information provided by natural language descriptions by maximizing the correlation between the text embedding and the image pixel embedding of the ground-truth class of the pixel. We use the official checkpoint  [^2] https://github.com/isl-org/lang-seg of ViT-L/16 that is trained on a mixture of seven datasets [44].
CLIP   [65]. CLIP aligns visual and textual representations in a shared embedding space through contrastive learning by maximizing the similarity between the embeddings of corresponding image-caption pairs while minimizing the similarity of non-matching pairs. CLIP was trained on a large and diverse dataset of image-caption pairs sourced from the internet including over 400 million image-text pairs. We use the official implementation and checkpoint  [^3] https://github.com/openai/CLIP with a ViT-L/14 as the backbone for our evaluations.
StableDiffusion (SD)   [70]. SD is a diffusion-based model used for generating high-quality images from text prompts. The model is trained to gradually remove noise from images, transforming random noise into coherent images that match the provided text descriptions. This model is trained on LAION5B [73] which contains over five billion of images paired with detailed captions. We follow DIFT [77]  [^4] https://github.com/Tsingularity/dift to extract features from SD and we use the checkpoint SD2.1 for our evaluation. We use the features from block index 1 for all tasks. The noise timestep is set to 100 by default. We use null-prompt as the text condition.
StableVideoDiffusion (SVD)   [12]. SVD is an extension of SD from image generation to video generation by incorporating additional temporal modules. SVD is first initialized from an image-level pretrained diffusion checkpoint (SD2.1), then is further finetuned on 10 million videos. We use their publicly released image-to-video variant (SVD-xt)  [^5] https://huggingface.co/stabilityai/stable-video-diffusion-img2vid-xt. We build our feature extractor pipeline following DIFT [77] and extract the features from index 1 for all tasks. The noise timestep is set to 25 by default. We use the first-frame image as the condition for all the cross-attention modules while we use the unconditional version for the latent input of the UNet – we concatenate an all-zero vector to the framewise embeddings. Each time, we feed a video clip of 25 frames to SVD to process the features.
V-JEPA   [11]. V-JEPA aims to learn robust visual representations by predicting future states of visual data. This model is pretrained on a mixed video dataset containing over 2 million of videos. We take their official implementation  [^6] https://github.com/facebookresearch/jepa and the checkpoint ViT-L/16 under resolution of 224$\times$224. We obtain their per-patch representation by removing the last pooling and linear layers. Each time, we feed a video clip of 16 frames to V-JEPA to process the features.
Swin3D   [93]. Swin3D adapts the Swin Transformer to handle 3D data, such as point clouds and volumetric data. We use the official checkpoint  [^7] https://github.com/microsoft/Swin3D that takes Swin3D-L as the backbone and is pretrained on the Structure3D dataset [110] with semantic segmentation as the target.
Datasets.   We evaluate the performance on two challenging indoor 3D VQA datasets: ScanQA [5] and SQA3D [52]. SQA3D features over 33K QA pairs, while ScanQA consists of more than 41K pairs. Each entry in these datasets includes a complex 3D indoor scene, a question, and corresponding answers. We use the splits provided by the respective datasets.
Optimization.   We keep the LLM parameters frozen and finetune the shallow visual projection Q-Former module [46] to align features from different encoders to the LLM input space. Different from [35], we train the Q-Former module from scratch for a fair comparison of all encoders. Following the approach of 3D-LLM, we pretrain the module for 10 epochs using 3D-Language dataset [35] and then finetune it on the training split of the two evaluation datasets for 35 epochs. Both stages use the AdamW [50] optimizer with a linear warmup and cosine decay learning rate scheduler. While longer training can further improve performance, trends stabilize after 35 training epochs.
A.3 Additional Evaluation Details for Registration
Dataset generation.   When generating corresponding partial scene point clouds from ScanNet dataset, due to memory constraint, we downsample the partial scene point clouds to 4,096 points each with the farthest point sampling (FPS) algorithm, if the number of points in $P_{1}$ and $P_{2}$ is over 4,096. We follow the same train/val split on the semantic segmentation task in our partial scene registration task.
Optimization.   We follow REGTR [95] to adopt a Transformer cross-encoder module to enable cross-reasoning of the foundation model features from two point clouds, followed by a lightweight decoder to obtain the corresponding position of each point in the other point cloud, forming altogether $N_{1}+N_{2}$ pairs of correspondences, where $N_{1}$ and $N_{2}$ are the number of points in $P_{1}$ and $P_{2}$, respectively. Afterward, the rotation $\mathbf{R}$ and the translation $\mathbf{t}$ can be obtained in a closed-form solution solved by a weighted version of the Kabsch-Umeyama [39, 81] algorithm. We use Adam [41] for optimization and train our model for 30 epochs.
A.4 License of Dataset Used
In this section, we list the licenses of all the datasets we have used during our evaluation:
  • ScanNet [20]: MIT License.
  • ScanQA [5]: CC BY-NC-SA 3.0 License.
  • SQA3D [52]: CC-BY-4.0 License.
  • ScanRefer [16]: CC BY-NC-SA 3.0 License.
  • 3DLanguage-Data [35]: MIT License.
Appendix B Additional Experiment Results
B.1 Evaluation Curves during Different Training Stages
We show the evaluation curves for the partial scene registration in Figure A. We can observe that the performance ranking of different foundation models stays mainly unchanged throughout the training process.
Stable Diffusion BLEU-1 BLEU-4 METEOR ROUGE CIDEr
@@ B.1 Evaluation Curves during Different Training Stages
Table 7: Evaluation of diffusion noise level and feature layers when using StableDiffusion [70] for feature extraction. The setting we choose are highlighted in bold.
B.2 Diffusion Noise Level and Feature Layer
@@
B.3 Additional Qualitative Results

### v2

@@ 3.2 Vision-Language Reasoning
Table 2: Evaluation of vision-language reasoning on ScanQA [5] and SQA3D [55] datasets. The top-2 results for each metric are shown in red and green, respectively. The 3D-LLM results [36] are shown for reference, indicating the relative position of our evaluation results with respect to the leading models trained on this task.
Figure 4: Evaluation curves on the ScanQA benchmark. The $x$-axis demonstrates models trained for different epochs. DINOv2 exhibits clearly superior performance.
@@ 3.3 Visual Grounding
Dataset.   We evaluate on the ScanRefer [16] dataset, which provides 51K text descriptions of 11K objects in 800 ScanNet scenes [20]. We report accuracy for unique, multiple, and overall categories, with unique referring to instances that have a unique semantic class in a given scene (easier).
@@ 3.4 Semantic Segmentation
Table 4: Evaluation of semantic segmentation on ScanNet [20] benchmark.
@@ 3.5 Registration: Geometric Correspondence
Model RR@0.05m (%) $\uparrow$ RR@0.1m (%) $\uparrow$ RR@0.2m (%) $\uparrow$ RRE () $\downarrow$ RTE (m) $\downarrow$
@@
To evaluate the geometric information contained in the VFM features, we design the following new task, partial scene registration, based on the point cloud registration [99, 51] task that performs homography estimation between two point clouds. From a complete point cloud representing the entire scene, we sample two point clouds $P_{1}\in\mathbb{R}^{N_{1}\times 3}$ and $P_{2}\in\mathbb{R}^{N_{2}\times 3}$ within the scene, corresponding to two sets of consecutive viewpoints which have a certain amount of overlapped region but are displaced with a homography transformation. Our goal is to find the homography matrix $H$ that correctly transforms the points in $P_{1}$ to register with $P_{2}$. Compared to the semantic segmentation task evaluated in Section 3.4, the partial scene registration task requires the foundation model features to have the capability of finding geometric correspondence for registration, which cannot be achieved simply by finding the correspondence according to semantic understanding. For example, in semantic correspondence, we may find two semantically similar points, one on the left side of the sofa in $P_{1}$, while the other on the right side of the sofa in $P_{2}$. As a result, if we register the two partial point clouds solely based on semantic correspondence, we will fail to find the correct homography to align one point cloud with the other. The VFMs need to be equipped with geometric understanding capability to achieve decent performance on our partial scene registration task.
Dataset. We build our partial scene registration benchmark based on ScanNet [20] dataset. For each scene in ScanNet, we choose views #0 $\sim$ #31 and views #32 $\sim$ #63 to render $P_{1}$ and $P_{2}$, respectively, so that they can have a certain level of overlap that allows the registration of two partial point clouds. Afterwards, $P_{2}$ is transformed by a homography $H$ that consists of a rotation $\mathbf{R}\in\mathrm{SO}(3)$ and a translation $\mathbf{t}\in\mathbb{R}^{3}$. $\mathbf{R}$ is created by a randomly generated quaternion $\mathbf{q}\in\mathbb{R}^{4}$ for each scene, while each component of $\mathbf{t}$ is randomly sampled from the uniform distribution $[-1.0\mathrm{m},1.0\mathrm{m}]$.
Optimization.   We follow REGTR [99] to adopt a transformer cross-encoder module to enable cross-reasoning of the foundation model features from two point clouds, followed by a lightweight decoder to obtain the corresponding position of every point in the other point cloud for all the $N_{1}+N_{2}$ points in both point clouds, forming altogether $N_{1}+N_{2}$ pairs of correspondences, where $N_{1}$ and $N_{2}$ are the number of points in $P_{1}$ and $P_{2}$, respectively. Afterward, the rotation $\mathbf{R}$ and the translation $\mathbf{t}$ can be obtained in a closed-form solution solved by a weighted version of the Kabsch-Umeyama [40, 84] algorithm. We use Adam [42] for optimization and train our model for 30 epochs, and follow REGTR to adopt Registration Recall (RR), Relative Rotation Error (RRE), and Relative Translation Error (RTE) as evaluation metrics.
@@ 4 Analysis
The purpose of this section is to provide additional exploration towards the optimal usage of visual foundation models. The selection of encoding methods requires consideration of the trade-off between memory usage, running time, and performance. We will dive into complexity analysis and the study of design choices for various and a combination of foundation models. More visualization, ablation experiments, and elaboration on the limitations, broader impact, and future direction are presented in the supplementary material.
@@ 4.1 Complexity Analysis
We compare memory usage, computation time, and model performance (vision-language reasoning on ScanQA) in Table 6 and Figure 6. Our findings show that image encoders generally require less time to process a sample compared to video and 3D encoders. And diffusion-based models, when used for feature extraction, require significantly more memory than other discriminative models. However, the drawbacks in running time become evident for 2D backbones, especially image encoders, when attempting to obtain a scene embedding by aggregating multi-view image embeddings. To illustrate this, we consider a 300-frame video as an exemplar of posed 2D information for a complex scene (a 10-second video at 30 FPS). As the length of the video increases, 2D methods, which necessitate feature extraction for each image frame, rapidly consume a substantial amount of time to process a single scene. In contrast, a 3D point encoder requires significantly less time to process a scene. Nevertheless, 3D encoders exhibit relatively poor model performance, which can be attributed to the scarcity of training data. To fully demonstrate their potential in scene understanding tasks, efforts should be directed toward enhancing the generalizability of 3D foundation models. All analyses and computations are performed on an NVIDIA A100 GPU.
@@ 4.2 Ablation Study – Insights into Optimal Usage of Visual Foundation Models
Combination of multiple encoders.   We explore whether a mixture of foundation models (experts) has the potential to strengthen the capability of 3D scene understanding. We experiment on the 3D semantic segmentation task with three feature sources: LSeg, StableDiffusion, and Swin3D. When combining different feature sources, we concatenate all features along the channel dimension for every point in the point cloud. The results are shown in Figure 8. After combining features from different sources, there exists a potential that the semantic understanding capability can be boosted in a mixture of experts manner. However, it is not necessarily true that combining the best features will lead to the best performance. For example, LSeg (1) has stronger capability on semantic segmentation than StableDiffusion (2) and Swin3D (3) individually, but it is StableDiffusion + Swin3D (2+3) that reaches the best performance when combining two features together.
[…]
@@
Table 7: Evaluation of diffusion noise level and feature layers when using StableDiffusion [73] for feature extraction. The settings we choose are highlighted with the grey color.
4.3 Diffusion Noise Level and Feature Layer
@@
5 Conclusion
This paper presents the first comprehensive analysis of leveraging visual foundation models for complex 3D scene understanding. We explore the strengths and weaknesses of models designed for various modalities and trained with different objectives. Our study reveals the superior performance of DINOv2, the advantages of video models in object-level tasks, and the benefits of diffusion models in geometric registration tasks. Surprisingly, we find limitations of language-pretrained models in language-related tasks. The extensive analysis suggests that a more flexible encoder selection and fusion can play a crucial role in future scene understanding and multimodal reasoning tasks.
Supplementary Material
Appendix A Additional Experiment Details
In this section, we provide a detailed introduction of all the visual foundation models we have evaluated, including the checkpoints we use and how we extract feature representations from the encoder backbones.
A.1 Evaluated Visual Foundation Models
Our evaluation and analysis are conducted mainly on the seven models listed in Table 1 in the main paper. We have chosen models such that they cover most of the backbones used by recent 3D scene understanding and reasoning work. In this part, we discuss all the models we have used in our experiments and explain their pretraining objective, the dataset used for pretraining, the public checkpoints we choose, and the method we leverage to extract features from their backbones. We start with image foundation models, and then video and 3D models.
DINOv2 [61].   DINOv2 leverages an image-wise contrastive objective by minimizing the distance of features from the same samples and maximizing those from different samples. It also includes a patch-wise denoising objective by performing a reconstruction from masked inputs. It is trained on a large-scale image dataset, LVD-142M [61], which contains 142 million unlabeled images. We take the standard DINOv2 implementation [^1] https://github.com/facebookresearch/dinov2 and use the pretrained ViT-L/14 checkpoint for our evaluations.
LSeg [46].   LSeg aims to align visual features from images with the corresponding semantic information provided by natural language descriptions by maximizing the correlation between the text embedding and the image pixel embedding of the ground truth class of the pixel. We use the official checkpoint [^2] https://github.com/isl-org/lang-seg of ViT-L/16 that is trained on a mixture of seven datasets [46].
CLIP[68].    CLIP aligns visual and textual representations in a shared embedding space through contrastive learning by maximizing the similarity between the embeddings of corresponding image-caption pairs while minimizing the similarity of non-matching pairs. CLIP is trained on a large and diverse dataset of image-caption pairs sourced from the Internet including over 400 million image-text pairs. We use the official implementation and checkpoint [^3] https://github.com/openai/CLIP with ViT-L/14 as the backbone for our evaluation.
StableDiffusion (SD) [73].   SD is a diffusion-based model used for generating high-quality images from text prompts. The model is trained to gradually remove noise from images, transforming random noise into coherent images that match the provided text descriptions. It is trained on LAION5B [76] which contains over five billion of images paired with detailed captions. We follow DIFT [80] [^4] https://github.com/Tsingularity/dift to extract features from SD and we use the SD2.1 checkpoint for our evaluation. We use the features from block index 1 for all tasks. The noise timestep is set to 100 by default. We use null-prompt as the text condition.
StableVideoDiffusion (SVD) [12].   SVD is an extension of SD from image generation to video generation by incorporating additional temporal modules. SVD is first initialized from an image-level pretrained diffusion checkpoint (SD2.1), and then further finetuned on 10 million videos. We use their publicly released image-to-video variant (SVD-xt)  [^5] https://huggingface.co/stabilityai/stable-video-diffusion-img2vid-xt. We build our feature extractor pipeline following DIFT [80] and extract the features from block index 1 for all tasks. The noise timestep is set to 25 by default. We use the first-frame image as the condition for all the cross-attention modules, while we use the unconditional version for the latent input of the UNet – we concatenate an all-zero vector to the framewise embeddings. Each time, we feed a 25-frame video clip to SVD to process the features.
V-JEPA [11].   V-JEPA aims to learn robust visual representations by predicting future states of visual data. This model is pretrained on a mixed video dataset containing more than 2 million videos. We take their official implementation [^6] https://github.com/facebookresearch/jepa and the ViT-L/16 checkpoint with a resolution of 224$\times$224. We obtain their per-patch representation by removing the last pooling and linear layers. Each time, we feed a 16-frame video clip to V-JEPA to process the features.
Swin3D [97].   Swin3D adapts the Swin Transformer to handle 3D data, such as point clouds and volumetric data. We use the official checkpoint [^7] https://github.com/microsoft/Swin3D that takes Swin3D-L as the backbone and is pretrained using the Structure3D dataset [114] with semantic segmentation as the target.
Datasets.   We evaluate the performance on two challenging indoor 3D VQA datasets: ScanQA [5] and SQA3D [55]. SQA3D features over 33K QA pairs, while ScanQA consists of more than 41K pairs. Each entry in these datasets includes a complex 3D indoor scene, a question, and the corresponding answers. We use the splits provided by the respective datasets.
Optimization.   We keep the LLM parameters frozen and finetune the shallow visual projection Q-Former module [48] to align the features of different encoders with the LLM input space. Unlike 3D-LLM [36], we train the Q-Former module from scratch for a fair comparison of all encoders. Following the approach of 3D-LLM, we pretrain the module for 10 epochs using the 3D-Language dataset [36] and then finetune it on the training split of the two evaluation datasets for 35 epochs. Both stages use the AdamW [53] optimizer with a linear warm-up and cosine decay learning rate scheduler. Although longer training can further improve performance, trends stabilize after 35 training epochs.
A.3 Additional Evaluation Details for Registration
Dataset generation.   When generating the corresponding partial scene point clouds from the ScanNet dataset, due to memory constraint, we downsample the partial scene point clouds to 4,096 points each with the farthest point sampling (FPS) algorithm, if the number of points in $P_{1}$ and $P_{2}$ is greater than 4,096. We follow the same train/val split on the semantic segmentation task in our partial scene registration task.
A.4 License of Datasets Used
We list the licenses of all the datasets we have used during our evaluation:
  • ScanNet [20]: MIT License.
  • ScanQA [5]: CC BY-NC-SA 3.0 License.
  • SQA3D [55]: CC-BY-4.0 License.
  • ScanRefer [16]: CC BY-NC-SA 3.0 License.
  • 3D-Language-Data [36]: MIT License.
Appendix B Additional Experimental Results
B.1 Comparison Between Scene-level and Object-centric Models
Uni3D [116] is a general transformer-based 3D foundation model pretrained on a mixture of four object-centric datasets [52]. Focusing on object-centric understanding, it has a restriction on the number of input points and output dimensions. In contrast, Swin3D [97] is pretrained on a smaller scene-level dataset [113], but is designed to focus more on understanding scene-level information. To demonstrate the performance of Uni3D [^8] https://github.com/baaivision/Uni3D, we conduct experiments with its features on our evaluation benchmarks. More specifically, following the part segmentation details in Uni3D’s appendix (Sec. B), we use Uni3D-giant, selecting features from the 16th, 28th, and 40th (last) layers to form grouped point patches. We then employ PointNet++’s [91] feature propagation to up-sample group features into point-wise features. It is worth noting that Uni3D’s ScanNet visualizations in their paper were achieved by applying Uni3D to each object instance based on ground truth instance segmentation, not by direct application to the whole scene.
Model VQA (CIDEr) $\uparrow$ Grounding (Acc.) $\uparrow$ Segmentation (mIoU) $\uparrow$ Registration (RTE) $\downarrow$
Swin3D 70.9 51.6 35.2 0.23
Uni3D 63.1 51.1 2.7 0.08
Table A: Comparison between Uni3D and Swin3D on four of our evaluation tasks. Object-centric and scene-centric methods demonstrate significant differences.
The results are shown in Table A. From the table we have several interesting findings:
  • For scene-level tasks (3D VQA and Semantic Segmentation): Uni3D underperforms the scene-level pretrained Swin3D model. This is likely due to the object-centric pretraining recipe of Uni3D, causing the failure of feature extraction on large scenes with orders of magnitude more points than single objects.
  • For object-centric tasks (3D object grounding): Uni3D achieves comparable results with Swin3D. However, some grounding questions require not only object-level semantics, but also inter-object relationship and global room information, which Uni3D lacks. We believe that combining object-centric and scene-level representations would be an impactful future direction to achieve better object grounding in complex 3D scenes.
  • For geometric understanding task (Registration): Uni3D achieves better performance than Swin3D, suggesting that geometric knowledge from object-centric pretraining generalizes well to scene-level geometric matching, especially given the task’s use of downsampled partial scenes bridging the distribution gap between object-level and scene-level point clouds.
B.2 Comparison Between Semantic and Instance Segmentation Models
Segment Anthing Model (SAM) [43] is an open-world instance segmentation model pretrained on a very large dataset SA-1B [43]. In Table B, we compare the performance of SAM with LSeg [46], a semantic segmentation model. For SAM, we use the official pretrained model checkpoint with ViT-L as the backbone encoder, matching the model size with other visual foundation models in our experiments.
Model VQA (CIDEr) $\uparrow$ Grounding (Acc.) $\uparrow$ Segmentation (mIoU) $\uparrow$ Registration (RTE) $\downarrow$
LSeg 71.0 50.4 47.5 0.59
SAM 68.6 50.1 30.9 0.09
Table B: Comparison between SAM and LSeg on four of our evaluation tasks. Instance-aware segmentation and semantic-aware segmentation methods demonstrate significant differences.
With the results in Table B, we offer the following analysis:
  • First, it is crucial to highlight the fundamental differences between LSeg and SAM. LSeg is designed to perform language-driven semantic image segmentation, providing semantic-aware representations. In contrast, SAM is primarily an instance segmentation model that focuses on local representations and excels in detecting edges. These distinctions result in varied performance on the four tasks in our evaluation.
  • Among the four tasks, 3D VQA and semantic segmentation require a deep semantic understanding of the 3D scenes, where LSeg naturally outperforms SAM. For 3D grounding, both semantic and spatial understanding is necessary; therefore, LSeg and SAM exhibit similar performance in this task. The registration task, however, demands matching point clouds using distinguishable local features. Here, SAM’s ability to provide precise local features positions it as a strong performer in this geometry-oriented task
  • Overall, SAM is not well-suited for numerous downstream tasks, particularly those requiring semantic comprehension. This conclusion is consistent with the previous study [70]. However, we also reveal that it excels in tasks that benefit from robust local feature representation.
B.3 Evaluation Curves During Different Training Stages
We show the evaluation curves for the partial scene registration task in Figure A. We can observe that the performance ranking of different foundation models stays mainly unchanged throughout the training process.
B.4 Additional Qualitative Results